OpenStudy (anonymous):

Find the center and the radius of the following circle: x^2 + y^2 - 6x + y = 3 Best explanation/answer earns medal!

OpenStudy (welshfella):

we can do this by completing the square to convert the equation to the form (x - a)^2 + ( y - b)^2 = r^2 Are you familiar with completing the square?

OpenStudy (anonymous):

i think so, i have the answers in my notes, but i'm just trying to remember what i did to get that answer

OpenStudy (anonymous):

Right now i have (x-3)^2 + (y + .5)^2

OpenStudy (welshfella):

x^2 - 6x = (x - 3)^2 - 9

OpenStudy (anonymous):

what about the y value?

OpenStudy (welshfella):

y^2 + y = (y + 0.5)^2 - 0.25

OpenStudy (anonymous):

so the center of the circle is (3,-.5) how do i find the radius?

OpenStudy (welshfella):

(x - 3)^2 + (y + 0.5)^2 = 3 + 9 + 0.25 = 12 .25 r^2 = 12.25 r = ?

OpenStudy (anonymous):

3.5??

OpenStudy (welshfella):

yep

OpenStudy (anonymous):

oh okay cool!

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

thanks!

OpenStudy (welshfella):

yw